# appendicesapc Appendix Table of Contents

### v1

@@ Abstract
Aligning large language models with human preferences improves interaction quality and safety by ensuring outputs better reflect human values. A promising strategy involves Reinforcement Learning from Human Feedback (RLHF), starting with collecting and ranking responses generated by a supervised fine-tuning model to refine alignment. Current methods (DPO) focus on learning from pairwise preference data, categorizing responses into preferred and less preferred pairs, and optimizing by maximizing pairwise margins. Recent studies have uncovered a substantial discrepancy between the theoretical aspirations of preference learning and its real-world results. Current preference alignment techniques underperform expectations, with ranking accuracies below $60\%$ on standard datasets. This suggests existing methods inadequately capture ideal preference relationships within sequences. To address this challenge, this paper introduces Direct Ranking Preference Optimization (DRPO), a novel method that views human preference alignment as a Learning-to-Rank (LTR) task. DRPO leverages NDCG, a widely used LTR metric, to optimize the ranking of responses within lists based on preference data, thereby enhancing ranking accuracies. Due to the nondifferentiability of NDCG, we propose diffNDCG loss, a differentiable approximation facilitated by a sorting network to simulate NDCG. Furthermore, to improve the quality of generated response, we propose a novel margin-based Adaptive Rank Policy Score. Extensive experiments have shown that DRPO outperforms existing baseline methods, enhancing the quality of the generated responses. The code is publicly available  [^1] Code and models can be found at https://anonymous.4open.science/r/drpo-align-2758.
1 Introduction
@@
Large language models (LLMs), trained on extensive and diverse datasets, can be prompted to demonstrate impressive capabilities across a broad range of tasks (Huang et al., 2024; Chiang et al., 2023; OpenAI et al., 2024; Touvron et al., 2023). However, due to the varied nature of their training data, these models sometimes produce content that may not align with human preferences, including fabricated answers, offensive comments, or harmful responses (Bai et al., 2022; Wang et al., 2023). To ensure the development of AI systems that are safe and controllable, this paper investigates learning tasks for LLMs that guide them to generate responses in alignment with human preferences.
@@
Human preference alignment has become an active research area. Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) is the first proposed method in this area. However, the optimization process of RLHF is complex, and its implementation introduces challenges due to unstable and costly training. Recent studies (Hong et al., 2024; Ethayarajh et al., 2024) have started to adopt alternatives to RLHF. For example, Direct Preference Optimization (DPO) (Rafailov et al., 2023) enables the extraction of the corresponding optimal policy in a closed form and derives a pairwise logistic loss directly from pairwise preference data. DPO eliminates the need for explicit reward modeling or reinforcement learning, thereby reducing the training costs associated with RLHF.
@@
Although significant progress has been made in human preference alignment, most existing methods primarily focus on pairwise human preferences, which involve evaluating human preferences by comparing preferred and less-preferred responses. Nevertheless, human preferences are not solely expressed as preferences and less preferences; they also manifest as ranking information, an aspect that has rarely been explored in previous research. In practice, ranking preference data is widely utilized. For instance, in the Ultrafeedback and VLfeedback datasets (Li et al., 2023a; Cui et al., 2023), multiple responses are generated using a Supervised Fine-tuning (SFT) model. These responses are then evaluated and ranked by leveraging advanced AI technologies such as GPT-4 (OpenAI et al., 2024). Moreover, as demonstrated by (Liu et al., 2024), presenting ranking preference data effectively distribute the costs associated with processing the prompt (Liu et al., 2023a).
@@
In this work, we approach human preference alignment as a listwise ranking problem, aiming to align LLMs by utilizing ranking preference data. The straightforward solutions for tackling this issue are to extend existing alignment methods from the pairwise scenarios to the ranking list scenarios. However, a major limitation is that they fail to utilize the relative strengths of the ranking preferences (Zhu et al., 2024). To address this, some studies leveraging ranking preference data have been proposed, proving their effectiveness for preference alignment (Yuan et al., 2023; Liu et al., 2024; Chen et al., 2024; Zhu et al., 2024; Choi et al., 2024). These methods predominantly achieve listwise preference alignment through two main strategies: either by utilizing the Plackett-Luce preference model, or by employing pairwise methods that incorporate listwise-aware weighting schemes (Liu et al., 2024).
@@
However, a mismatch persists between evaluation metrics and optimization objectives. LLMs’ performance is typically assessed using win rates, which measure how often one model’s responses are preferred over another’s, yet current alignment methods do not directly optimize this criterion. This mismatch means that optimizing current loss functions (Liu et al., 2024; Song et al., 2024) may not necessarily lead to higher win rates or improved human preference satisfaction (Chen et al., 2024), substantially hindering model preference alignment. In contrast, Pobrotyn and Bialobrzeski (2021); Qin et al. (2010a) have demonstrated that direct optimization of evaluation metrics is highly effective in traditional learning-to-rank tasks. Moreover, recent studies (Chen et al., 2024) have demonstrated a correlation between win rates and ranking accuracy, while revealing that current alignment methods achieve low ranking accuracy on standard preference datasets, but the relationship between win rates and more comprehensive metrics like NDCG remains unexplored. Therefore, aligning LLMs with human preferences by directly optimizing evaluation criteria and improving ranking accuracy remains a challenging task that has been scarcely explored to date.
@@
To tackle this challenge, we propose a novel method: Direct Ranking Preference Optimization (DRPO), which is a pioneering exploration to align LLMs with the ranking metric Normalized Discounted Cumulative Gain (NDCG), a metric from Learning to Rank (LTR) that accurately quantifies ranking accuracy and quality. Specifically, We introduce the Adaptive Rank Policy Score, a novel ranking strategy that maximizes the absolute likelihood of preferred responses while dynamically adjusting score margins between preferred and non-preferred responses based on their relative positions in the ranked list. Furthermore, we implement differentiable sorting networks to sort responses based on computed scores. This sorting method combines implementation simplicity and computational efficiency, while yielding doubly stochastic permutation matrices that preserve probability distributions and enable efficient differentiable optimization (Petersen et al., 2021). Additionally, due to the nondifferentiability of NDCG, we leverage permutation matrices develop a differentiable version of NDCG (i.e., diffNDCG) to serve as the loss function, simulating the NDCG metric. Optimizing the diffNDCG loss enhances model performance by prioritizing top-ranked responses and imposing stricter penalties for misplacing highly relevant items, without additional computational overhead compared to existing methods. The main contributions of our work are summarized as follows:
@@
$\bullet$ A novel method Direct Ranking Preference Optimization (DRPO) is developed to explore human preference alignment with ranking preference data. To our best knowledge, this paper is a pioneer exploration to align LLMs with ranking human preference data.
@@
$\bullet$ A novel ranking score computation strategy, the Adaptive Rank Policy Score, has been introduced to replace the classical computation method used in RLHF (Rafailov et al., 2023). It maximizes preferred response likelihood while dynamically adjusting score margins based on relative positions.
@@
$\bullet$ A novel differentiable NDCG (diffNDCG) metric has been developed to emulate the NDCG metric used in LTR. By optimizing diffNDCG, we can prioritize responses at the top of the ranking list and impose stricter penalties for inaccurately placing a top-ranked response in a lower position.
Figure 1: A diagram illustrating the two key steps of our method: (1) collecting ranking preference data, and (2) optimizing the policy with the collected ranking preference data. In Step 2, the Policy Advantage Score computes the score for each input response. Subsequently, the Differentiable Sorting Network sorts these responses based on their scores. We then compute the diffNDCG between the predicted scores and the ground-truth scores, and optimize the policy using the diffNDCG loss.
2 Preliminaries
@@
Prompt, Response and Policy. Let $\mathcal{X}$ and $\mathcal{Y}$ denote the set of prompts and the set of responses (action space), respectively. We use $x\in\mathcal{X}$ to represent a prompt, and $y\in\mathcal{Y}$ to represent a response. Given a prompt $x$, a large language model (LLM) generates a corresponding response $y$. This response $y$ is produced according to a policy $\pi_{\theta}(\cdot|x)$, which is a discrete distribution over $\mathcal{Y}$. We also define $\pi_{\text{ref}}(\cdot|x)$ as a discrete distribution over $\mathcal{Y}$, serving as the reference policy. The reference policy $\pi_{\rm ref}$ is derived from the Supervised Fine-tuning (SFT) model (Rafailov et al., 2023).
@@
Ranking Preference Data. The training dataset $D=\{x^{i},\mathbf{y}^{i},\mathbf{s}^{i}\}_{i=1}^{N}$ is composed of three elements: $x^{i}$ represents the $i$-th prompt; $\mathbf{y}^{i}=(y^{i}_{1},\ldots,y^{i}_{K})$ consists of a list of $K$ responses, typically generated by the SFT model; and $\mathbf{s}^{i}=(s^{i}_{1},\ldots,s^{i}_{K})\in[0,1]^{K}$ denotes the relevance scores of the responses $\mathbf{y}^{i}$ in relation to the prompt $x^{i}$. The relevance score $s_{j}^{i}$, generally obtained from AI (Huang et al., 2024; Jiang et al., 2023; Bai et al., 2023a; Chiang et al., 2023) and human feedback or a reward model, reflects how well the response $y_{j}^{i}$ corresponds to the prompt $x^{i}$. If the response $y^{i}_{j}$ is scored higher than $y^{i}_{l}$, it implies that $y^{i}_{j}$ is more closely aligned with human preferences compared to $y^{i}_{l}$.
@@
Alignment with Human Preferences Using Ranking Preference Data. Aligning LLM with human preferences involves utilizing the dataset of human preferences (e.g., the training dataset $D$) to refine the policy $\pi_{\theta}(y|x)$. Substantial progress has been made toward achieving this goal, with most existing studies (Rafailov et al., 2023; Hong et al., 2024) leveraging pairwise preference data, represented as $D_{P}=\{x^{i},(y^{i}_{1},y^{i}_{2}),(s^{i}_{1},s^{i}_{2})\}_{i=1}^{N}$ (i.e., the case $K=2$ for training dataset $D$). Unlike existing methods, we treat preference alignment as a listwise ranking problem (i.e., $K\geq 2$), which allows for a more effective exploitation of the complex preference relationships embedded within sequence.
Learning-to-Rank Task. Given the ranking preference data, the human preference alignment can be regarded as the Learning-to-Rank (LTR) task (Liu et al., 2009; Yu et al., 2019; Cao et al., 2007). When a user enters a query $x_{q}$, the LTR algorithm needs to reliably rank multiple candidate documents (e.g., texts, images and web pages) to ensure that the most relevant information is retrieved first so that the user can quickly find the information they need. Let $\mathbf{M}$ be the LTR model. Given a query ${x}_{q}$ and documents $\mathbf{y}_{d}$, $\mathbf{M}$ predicts relevance scores $\mathbf{s}$, assigning higher scores to superior documents.
@@
where $\pi_{\theta}$ is the LLM policy and $\theta$ is the corresponding parameters. Let $\ell:(\mathbf{s},\hat{\mathbf{s}_{\theta}})\rightarrow\mathbb{R}$ be the loss function. We will learn the parameters $\theta$ based on the empirical risk minimization principle:
@@
3 Proposed Methodology
@@
In this section, we introduce the proposed method DRPO, which comprises three primary components: (1) ranking score computation; (2) differentiable responses ranking; (3) diffNDCG loss. The graphical illustration of our proposed method is depicted in Figure 1.
3.1 Ranking Score Computation
@@
Policy Reference Ratio. The fundamental criterion for computing the ranking score is that more preferred responses should receive higher scores. A commonly used strategy to compute the ranking scores is Policy Reference Ratio proposed by Rafailov et al. (2023): let $\mathbf{M_{prr}}$ be the policy reference ratio model of human preference alignment, which can be expressed as:
$\mathbf{M_{prr}}(x,\mathbf{y};\pi_{\theta})=\Big{(}\beta\log\frac{\pi_{\theta}(y_{1}\mid x)}{\pi_{\text{ref}}(y_{1}\mid x)},...,\beta\log\frac{\pi_{\theta}(y_{K}\mid x)}{\pi_{\text{ref}}(y_{K}\mid x)}\Big{)},$ (2)
@@
where $\beta$ is the hyper-parameter to control the KL divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$.
@@
Adaptive Rank Policy Score. While the Policy Reference Ratio defined in Eq. 2 has been widely adopted in various methods (Liu et al., 2024), it emphasizes the relative likelihood between the policy model $\pi_{\theta}$ and a reference model $\pi_{\mathrm{ref}}$ rather than directly maximizing the absolute likelihood of the preferred response. Consequently, a high Policy Reference Ratio score may coincide with low absolute likelihood for preferred responses (Meng et al., 2024), leading to sub-optimal performance in real-world generation tasks, where high absolute likelihoods are essential for producing high quality outputs (Holtzman et al., 2018; Fan et al., 2018). To address this, we focus on the log-likelihood of generated sequences and establish a length normalized basic scores function based on log-likelihood:
$\mathbf{s}(x,\mathbf{y};\pi_{\theta})=\Big{(}\frac{1}{|y_{1}|}\log\pi_{\theta}(y_{1}\mid x),...,\frac{1}{|y_{K}|}\log\pi_{\theta}(y_{K}\mid x)\Big{)},$ (3)
@@
Here, $|y|$ denotes the token length of $y$. This length normalization reduces bias towards shorter sequences (Yuan et al., 2023). Furthermore, when performing the Differentiable Swapping Operation (see Section 2), we calculate score differences between elements in the responses list. In these calculations, a common practice to enhance discrimination between high and low-quality responses is to incorporate a margin (Meng et al., 2024; Ethayarajh et al., 2024), ensuring preferred responses exceed dispreferred ones by at least a specified threshold. This margin-based methodes has been empirically demonstrated to enhance model generalization and improve the quality of generated responses (Touvron et al., 2023). To this end, we introduce a additional ranking-aware term $\gamma(y)$:
$\mathbf{s}(x,\mathbf{y};\pi_{\theta})=\Big{(}\frac{1}{|y_{1}|}\log\pi_{\theta}(y_{1}\mid x)+\gamma(y_{1}),...,\frac{1}{|y_{K}|}\log\pi_{\theta}(y_{K}\mid x)+\gamma(y_{K})\Big{)}.$ (4)
@@
The ranking-aware margin is then defined as the difference between $\gamma(y_{i})$ and $\gamma(y_{j})$ when comparing scores of two responses $y_{i}$ and $y_{j}$. This margin should effectively reflect quality differences among responses across the ranked list. Specifically, we assume adjacent responses have similar relevance, and design ranking-aware margin to satisfy three criteria: first, apply smaller margins for adjacent ranks, allowing fine-grained discrimination; second, assign larger margins for greater ranking disparities, emphasizing significant differences; finally, dynamically adjust margins based on relative score changes, maintaining discrimination across the quality levels while avoiding overemphasis on minor differences. Therefore, we define $\gamma(y)$ by combining a base weighted ranking position term with an exponential moving average estimate (Qin et al., 2010b) of past scores related to the ranking position:
$\gamma(y)$ $=\tau\cdot q(y)-\beta\cdot{V}_{q(y)},~{}\text{where}~{}{V}_{q(y)}$ $=\theta\cdot{V}_{q(y)}+(1-\theta)\cdot\frac{1}{|y|}\log\pi_{\theta}(y\mid x).$ (5)
@@
Here, $q(y)$ denotes the ranking position of response $y$ (e.g., 0 for the highest-ranked response), and $\tau$ is a positive constant factor representing the atomic margin between adjacent responses. ${V}_{q(y)}$ is the exponential moving average estimate of the log likelihood at rank $q(y)$ for dynamically tracking historical changes. $\theta\in[0,1]$ is the parameter controlling the update rate, while $\beta$ determines the influence of the historical estimate on the current score. Building upon adaptive ranking-aware term $\gamma(y)$, we propose the novel Adaptive Ranking Policy Score:
$\mathbf{M}_{\rm arp}(x,\mathbf{y};\pi_{\theta})$ $=\Big{(}s(x,y_{1};\pi_{\theta}),s(x,y_{2};\pi_{\theta}),...,s(x,y_{K};\pi_{\theta})\Big{)},$ (6)
@@
where $s(x,y;\pi_{\theta})=\log{\pi_{\theta}(y\mid x)}/{|y|}+\tau\cdot q(y)-\beta V_{q(y)}$. Detailed and rigorous experiments demonstrate that our Adaptive Ranking Policy Score consistently outperforms the Policy Reference Ratio score across a wide range of metrics and diverse datasets (see Section 4).
3.2 Differentiable Responses Ranking
Figure 2: Sort $(10,2,4,8)$ in descending order using a sorting network. (a) The structure of the sorting network; (b) the swapping operation in the first layer.

### v2

# Holistic Utility Preference Learning for Listwise Alignment

@@ Abstract
Aligning large language models with human preferences is essential for improving interaction quality and safety by ensuring outputs better reflect human values. A promising strategy involves Reinforcement Learning from Human Feedback (RLHF), starting with collecting and ranking responses generated by a supervised fine-tuning model to refine alignment. Existing methods such as Direct Preference Optimization (DPO) focus on pairwise comparisons, categorizing responses into preferred and less preferred pairs and optimizing pairwise margins. However, this pairwise approach cannot capture the holistic ranking relationships among multiple responses or effectively leverage the rich preference information available in list-wise comparisons. To address this challenge, this paper introduces Direct Ranking Preference Optimization (DRPO), a novel method that views human preference alignment as a Learning-to-Rank (LTR) task. Unlike pairwise methods, DRPO optimizes the preference ranking of entire response lists by computing holistic utility scores through NDCG, a standard LTR metric. To enable end-to-end optimization with the non-differentiable NDCG, we propose diffNDCG loss, a differentiable approximation facilitated by a sorting network. Furthermore, we introduce a novel margin-based Adaptive Rank Policy Score to enhance the discriminative quality of generated responses. Extensive experiments have shown that DRPO outperforms existing methods, enhancing the quality of the generated responses. The code is available here
I Introduction
@@
Large language models (LLMs), trained on extensive and diverse datasets, can be prompted to demonstrate impressive capabilities across a broad range of tasks [1, 2, 3, 4, 5, 6]. However, due to the varied nature of their training data, these models sometimes produce content that may not align with human preferences, including fabricated answers, offensive comments, or harmful responses [7, 8]. To ensure the development of AI systems that are safe and controllable, this paper investigates learning tasks for LLMs that guide them to generate responses in alignment with human preferences.
@@
Human preference alignment has emerged as an active research area, pioneered by Reinforcement Learning from Human Feedback (RLHF) [9]. However, the RLHF optimization process is complex and its implementation introduces challenges due to unstable and costly training. Recent studies [10, 11] have started to adopt alternatives to RLHF. For example, Direct Preference Optimization (DPO) [12] enables the extraction of the corresponding optimal policy in a closed form and derives a pairwise logistic loss directly from pairwise preference data. DPO eliminates the need for explicit reward modeling or reinforcement learning, thereby reducing the training costs associated with RLHF.
@@
Although significant progress has been made in human preference alignment, most existing methods primarily focus on pairwise human preferences, which involve evaluating human preferences by comparing preferred and less-preferred responses. Nevertheless, human preferences are not solely expressed as preferences and less preferences; they also manifest as ranking information, an aspect that has rarely been explored in previous research. In practice, ranking preference data is widely utilized. For instance, in the Ultrafeedback and VLfeedback datasets [13, 14], multiple responses are generated using a Supervised Fine-tuning (SFT) model. These responses are then evaluated and ranked by leveraging advanced AI technologies such as GPT-4 [3]. Moreover, as demonstrated by [15], presenting ranking preference data effectively distribute the costs associated with processing the prompt [16].
@@
In this work, we approach human preference alignment as a listwise ranking problem. While straightforward solutions extend existing pairwise alignment methods to ranking scenarios, they fail to fully utilize the relative preference strengths within ranked lists [17]. Recent studies have attempted to address this limitation through two main strategies: utilizing the Plackett-Luce preference model or employing pairwise methods with listwise-aware weighting schemes [18, 19, 17, 20, 15].
@@
However, a mismatch persists between evaluation metrics and optimization objectives in all current methods. Alignment quality is typically assessed using win rates, which measure how often one model’s responses are preferred over another’s, yet existing alignment methods do not directly optimize this criterion. This mismatch means that optimizing current loss functions [15, 21] may not necessarily lead to higher win rates or improved human preference satisfaction [19]. While traditional learning-to-rank tasks have demonstrated that directly optimizing evaluation metrics is highly effective [22, 23], and recent studies [19] reveal correlations between win rates and ranking accuracy, current alignment methods achieve low ranking accuracy on standard datasets. The relationship between win rates and more comprehensive ranking metrics like NDCG remains unexplored, making direct optimization of evaluation criteria a critical yet challenging research gap.
@@
To tackle this challenge, we propose Direct Ranking Preference Optimization (DRPO), a novel method that directly aligns LLMs with Normalized Discounted Cumulative Gain (NDCG), a standard Learning-to-Rank metric that accurately quantifies ranking quality. Specifically, we introduce the Adaptive Rank Policy Score, a ranking strategy that maximizes the likelihood of preferred responses while dynamically adjusting score margins based on relative positions in the ranked list. Furthermore, we implement differentiable sorting networks to sort responses based on computed scores. This sorting method combines implementation simplicity and computational efficiency, while yielding doubly stochastic permutation matrices that preserve probability distributions and enable efficient differentiable optimization [24]. Additionally, due to the nondifferentiability of NDCG, we leverage permutation matrices develop a differentiable version of NDCG (i.e., diffNDCG) to serve as the loss function, simulating the NDCG metric. Optimizing the diffNDCG loss enhances model performance by prioritizing top-ranked responses and imposing stricter penalties for misplacing highly relevant items, without additional computational overhead compared to existing methods. The main contributions of our work are summarized as follows:
@@
$\bullet$ A novel method Direct Ranking Preference Optimization (DRPO) is developed to explore listwise human preference alignment with ranking preference data.
@@
$\bullet$ A novel ranking score computation strategy, the Adaptive Rank Policy Score, has been introduced, which maximizes the likelihood of preferred responses while dynamically adjusting score margins based on relative positions.
@@
$\bullet$ A novel differentiable NDCG (diffNDCG) metric has been developed to emulate the NDCG metric used in LTR. By optimizing diffNDCG, we prioritize top-ranked responses while penalizing their misplacement in lower positions.
II Related Works
Human Preference Alignment. Although LLMs demonstrate impressive capabilities, they sometimes produce harmful content, leading to safety concerns. To address this, Reinforcement Learning with Human Feedback (RLHF) [7, 25] aligns LLMs with human preferences by training a reward model using the Bradley-Terry model [26] and fine-tuning with reinforcement learning algorithms like PPO [27]. However, RLHF suffers from instability and high computational costs. To reduce reliance on reinforcement learning, Direct Preference Optimization (DPO) [12] derives a closed-form optimal policy with pairwise logistic loss, eliminating explicit reward modeling. RSO [16] employs rejection sampling for more accurate policy estimation and provides a unified framework extending SLiC [28] and DPO.
Most recent works [29, 30, 18, 31] focus on pairwise preferences, with limited attention to listwise settings. RRHF [18] uses pairwise hinge loss but treats pairs independently rather than modeling the complete ranking structure. PRO [21] employs the Plackett-Luce model [32] with list MLE loss. LiPO [15] frames preference alignment as learning-to-rank and uses DCG-weighted pairwise logistic loss (lambda loss) [33] to indirectly optimize NDCG. However, [22] showed that direct NDCG optimization yields superior performance in traditional learning-to-rank tasks.
Building upon this insight, we implement differentiable sorting networks to obtain differentiable permutation matrices of responses. It enable us to develop a differentiable NDCG(diffNDCG), allowing direct optimization of NDCG for ranked lists to more effectively align LLMs with human preferences.
Learning-to-Rank Task. Learning-to-Rank (LTR) task is a well-studied field with extensive literature, primarily due to its practical applications in web search and recommendation systems [34, 35, 36, 37, 38, 39, 40, 41]. Traditional research in Learning to Rank (LTR) has concentrated on developing robust ranking objectives, including pointwise, pairwise, and listwise approaches. RankSVM [42] and RankNet [43] utilize pairwise hinge loss and pairwise logistic loss, respectively, to optimize ranking performance. ListMLE and Softmax losses are two representative listwise losses introduced by [37]. LambdaRank [44] employs a pairwise logistic loss with lambda weights and achieves strong empirical performance improvements compared to RankNet. Furthermore, numerous methods have been proposed to directly optimize the non-smooth NDCG metric. For instance, SoftRank[45],ApproxNDCG [23] employs rank distributions to smooth NDCG. PiRank [46] and NeuralNDCG [22] approximate the non-continuous sorting operator using NeuralSort [47] to smooth NDCG. Inspired by these studies, we utilize differentiable sorting networks to smooth NDCG and introduce the diffNDCG loss function for preference alignment.
III Preliminaries
@@
Prompt, Response and Policy. Let $\mathcal{X}$ and $\mathcal{Y}$ denote the set of prompts and the set of responses (action space), respectively. We use $x\in\mathcal{X}$ to represent a prompt, and $y\in\mathcal{Y}$ to represent a response. Given a prompt $x$, a large language model (LLM) generates a corresponding response $y$. This response $y$ is produced according to a policy $\pi_{\theta}(\cdot|x)$, which is a discrete distribution over $\mathcal{Y}$. We also define $\pi_{\text{ref}}(\cdot|x)$ as a discrete distribution over $\mathcal{Y}$, serving as the reference policy. The reference policy $\pi_{\rm ref}$ is derived from the Supervised Fine-tuning (SFT) model [12].
@@
Ranking Preference Data. The training dataset $D=\{x^{i},\mathbf{y}^{i},\mathbf{s}^{i}\}_{i=1}^{N}$ is composed of three elements: $x^{i}$ represents the $i$-th prompt; $\mathbf{y}^{i}=(y^{i}_{1},\ldots,y^{i}_{K})$ consists of a list of $K$ responses, typically generated by the SFT model; and $\mathbf{s}^{i}=(s^{i}_{1},\ldots,s^{i}_{K})\in[0,1]^{K}$ denotes the relevance scores of the responses $\mathbf{y}^{i}$ in relation to the prompt $x^{i}$. The relevance score $s_{j}^{i}$, generally obtained from AI [1, 48, 49, 2] and human feedback or a reward model, reflects how well the response $y_{j}^{i}$ corresponds to the prompt $x^{i}$. If the response $y^{i}_{j}$ is scored higher than $y^{i}_{l}$, it implies that $y^{i}_{j}$ is more closely aligned with human preferences compared to $y^{i}_{l}$.
@@
Alignment with Human Preferences Using Ranking Preference Data. Aligning LLM with human preferences involves utilizing the dataset of human preferences (e.g., the training dataset $D$) to refine the policy $\pi_{\theta}(y|x)$. Substantial progress has been made toward achieving this goal, with most existing studies [12, 10] leveraging pairwise preference data, represented as $D_{P}=\{x^{i},(y^{i}_{1},y^{i}_{2}),(s^{i}_{1},s^{i}_{2})\}_{i=1}^{N}$ (i.e., the case $K=2$ for training dataset $D$). Unlike existing methods, we treat preference alignment as a listwise ranking problem (i.e., $K\geq 2$), which allows for a more effective exploitation of the complex preference relationships embedded within sequence.
Learning-to-Rank Task. When a user enters a query $x_{q}$, the LTR algorithm [34, 39, 37] aims to accurately rank candidate documents (including texts, images, and other modalities) by relevance, prioritizing the most pertinent information for efficient user retrieval. Let $\mathbf{M}$ denote the LTR model that assigns relevance scores $\mathbf{s}$ between a query ${x}_{q}$ and documents $\mathbf{y}_{d}$, where higher scores indicate more relevant documents.
@@
[…]
@@
Figure 1: Illustration of two key steps of our method: (1) collecting ranking preference data, and (2) optimizing the policy with the collected ranking preference data. In Step 2, the Adaptive Rank Policy Score computes the score for each input response. Subsequently, the Differentiable Sorting Network sorts these responses based on their scores. We then compute the diffNDCG between the predicted scores and the ground-truth scores, and optimize the policy using the diffNDCG loss.
IV Proposed Methodology
@@
In this section, we introduce the proposed method DRPO, which comprises three primary components: (1) ranking score computation; (2) differentiable responses ranking; (3) diffNDCG loss. Figure 1 shows the overview of our method.
IV-A Ranking Score Computation
@@
r Policy Reference Ratio. The fundamental criterion for computing the ranking score is that more preferred responses should receive higher scores. A commonly used strategy to compute the ranking scores is Policy Reference Ratio proposed by [12]: Let $\mathbf{M_{prr}}$ denote the policy reference ratio model for human preference alignment, which is defined as:
$\begin{gathered}\mathbf{M_{prr}}(x,\mathbf{y};\pi_{\theta})=\Big(\beta\log\frac{\pi_{\theta}(y_{1}\mid x)}{\pi_{\text{ref}}(y_{1}\mid x)},...,\\ \beta\log\frac{\pi_{\theta}(y_{K}\mid x)}{\pi_{\text{ref}}(y_{K}\mid x)}\Big),\end{gathered}$ (2)
@@
where $\beta$ denotes the regularization coefficient controlling the KL divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$.
@@
Adaptive Rank Policy Score. While the Policy Reference Ratio defined in Eq. 2 has been widely adopted in various methods [15], it emphasizes the relative likelihood between the policy model $\pi_{\theta}$ and a reference model $\pi_{\mathrm{ref}}$ rather than directly maximizing the absolute likelihood of the preferred response. Consequently, a high Policy Reference Ratio score may coincide with low absolute likelihood for preferred responses [50], leading to sub-optimal performance in real-world generation tasks, where high absolute likelihoods are essential for producing high quality outputs [51, 52]. To address this, we focus on the log-likelihood of generated sequences and establish a length normalized basic scores function based on log-likelihood:
$\begin{gathered}\mathbf{s}(x,\mathbf{y};\pi_{\theta})=\Big(\frac{1}{|y_{1}|}\log\pi_{\theta}(y_{1}\mid x),...,\\ \frac{1}{|y_{K}|}\log\pi_{\theta}(y_{K}\mid x)\Big),\end{gathered}$ (3)
@@
Here, $|y|$ denotes the token length of $y$. This length normalization reduces bias towards shorter sequences [18]. Furthermore, in the process of ranking responses based on their scores, pairwise score differences between elements are computed. A common practice to enhance discrimination between high and low-quality responses is to incorporate a margin [50, 11], ensuring preferred responses exceed dispreferred ones by at least a specified threshold. Empirical evidence suggests that margin-based approach improves both model generalization and response quality [4]. Therefore, we further introduce a ranking-aware term $\gamma(y)$.
$\begin{gathered}\mathbf{s}(x,\mathbf{y};\pi_{\theta})=\Big(\frac{1}{|y_{1}|}\log\pi_{\theta}(y_{1}\mid x)+\gamma(y_{1}),...,\\ \frac{1}{|y_{K}|}\log\pi_{\theta}(y_{K}\mid x)+\gamma(y_{K})\Big).\end{gathered}$ (4)
@@
The ranking-aware margin is then defined as the difference between $\gamma(y_{i})$ and $\gamma(y_{j})$ when comparing scores of two responses $y_{i}$ and $y_{j}$. This margin should effectively reflect quality differences among responses across the ranked list. Specifically, we assume adjacent responses have similar relevance, and design ranking-aware margin to satisfy three criteria: first, apply smaller margins for adjacent ranks, allowing fine-grained discrimination; second, assign larger margins for greater ranking disparities, emphasizing significant differences; finally, dynamically adjust margins based on relative score changes, maintaining discrimination across the quality levels while avoiding overemphasis on minor differences. Therefore, we define $\gamma(y)$ by combining a base weighted rank term with an exponential moving average estimate [53] of rank-related past scores:
$\gamma(y)$ $=\tau\cdot q(y)-\beta\cdot{V}_{q(y)},~\text{where}$~ (5)
${V}_{q(y)}$ $\leftarrow\theta\cdot{V}_{q(y)}+(1-\theta)\cdot\frac{1}{|y|}\log\pi_{\theta}(y\mid x).$
@@
Here, $q(y)$ is the rank of response $y$ (0 for the top-ranked response), and $\tau$ is a positive constant denoting the basic margin between adjacent responses. ${V}_{q(y)}$ is the exponential moving average estimate of the log likelihood at rank $q(y)$ for dynamically tracking historical changes. $\theta\in[0,1]$ is the parameter controlling the update rate, while $\beta$ determines the influence of the historical estimate on the current score. Based on the adaptive ranking-aware term $\gamma(y)$, we propose the novel adaptive ranking policy score:
$\mathbf{M}_{\rm arp}(x,\mathbf{y};\pi_{\theta})$ $=\Big(s(x,y_{1};\pi_{\theta}),.,s(x,y_{K};\pi_{\theta})\Big),$ (6)
@@
where $s(x,y;\pi_{\theta})=\log{\pi_{\theta}(y\mid x)}/{|y|}+\tau\cdot q(y)-\beta V_{q(y)}$. Detailed and rigorous experiments demonstrate that our Adaptive Ranking Policy Score consistently outperforms the Policy Reference Ratio score across a wide range of metrics and diverse datasets (see Section V).
IV-B Differentiable Responses Ranking